CH PAS Master Timeline

From New IAC Wiki
Jump to navigation Jump to search


This page will be used to document when each task is started/completed as a master reference for all work accomplished. The plan is to make my life easier when I have to look through months of work.



  • Attended intern orientation for first day at INL
  • Completed training this week as well as began NaI plateau setup


  • Put first coat of paint onto lead bricks to be used as shielding in PAS experiment


  • Moved steel table from LDS to clean room and began set up


  • Sheared a lead sheet into 6 rough 3"x3" squares to be used for attenuation measurements with NaI detectors
  • Lead sheet pieces were wiped down with a wet wipe, duct taped, rewiped, and hammered flat.


  • Began data collection for NaI plateau and operating point measurements
  • Began writing analysis code for NaI plateau measurements


  • Finished painting lead bricks and left to cure

Unknown time when tasks were completed

  • Gathered borated poly with Paul
  • Tested NaI detectors and found one working base



  • Moved dried/cured painted lead bricks from painting area to clean room
  • Cleaned each brick with a wet wipe before transition into clean room, will be allowed to air dry on steel table
    • 50 bricks total: 49 half bricks and 1 full brick
  • Steel table set up finished


  • Created initial lead brick tunnel


  • Finished collection and analysis of NaI detector plateau measurements, voltages ranging from 700V to 1400V
    • Analysis included Souce in & Source out plots, Signal:Noise, and background subtracted plots


  • Began data colleciton for He3 Borated Poly attenuation study


  • Finished data collection for He3 Borated Poly attenuation study
  • Began writing analysis code for Borated Poly attenuation study


  • Spent the day at INL with Chuting and Paul


  • Finished writing analysis code for He3 Borated Poly attenuation measurements and submitted plot to Paul
  • Began designing cradle for NaI detector using fusion360


  • Finished modeling cradle and sent designs to Jeff Burggraf who offered to print it for me


  • Began BorAl attenuation measurements including overnight background runs


  • First BorAl background run lasted 60000 seconds and ended at approximately 11:45am 6/26/21
  • Continuing to run Borated Poly attenuation measurements with BorAl surrounding He3 tube
  • Cradle for NaI detector has been printed and should be in lab Tuesday 6/29/21
  • Created github repository to advertise my poor coding practices
  • Began mapping shielding measurements for Lead, BorAl, and various Poly


  • Continued BorAl/Borated Poly attenuation measurements with He3 tube
  • Began working with MCNP
  • Continued to map out BorAl placement and shielding design for feasibility study
    • We need more BorAl


  • Finished BorAl/Borated Poly attenuation measurements
  • Got signal from HPGe detector
  • Discussed various poly box and shielding arrangements for feasibility study
    • Designs need to be drawn up
    • Lead tunnel needs to be made bigger to accommodate HPGe detector
      • Will probably need to model another cradle for HPGe
    • Need to source more regular poly and cut it to appropriate length and widths


  • Sourced regular poly and lead sheets, will be washed and brought to clean room today
  • Finalized design of poly box for feasibility study
    • Will have to cut some poly blocks to fit the needed dimensions
  • Finished designing poly brick thermalizing box and lead shielding
  • Brought poly down to LDS in dark room to be washed
  • Tony and Paul were able to get a good signal out of HPGe detector

Unknown time when tasks were completed

  • Found a fifth NaI detector (no base) with smaller crystal diameter (5.5cm)
  • Helped Tony and Joe moved DD generator from nuke eng department and subsequently lift into clean room
  • Moved HPGe detector into clean room and unpacked it, Paul has begun preliminary functionality testing



  • Installed cradle into lead tunnel
  • Discussed solid state physics with Dr. Forest regarding HPGe detector
  • Tested various arrangements to show that BorAl is most effective after borated poly and not before down beam axis
  • Received new daq from CoSE IT and Dr. Forest did preliminary setup


  • Finished and posted BorAl/BPoly/Poly attenuation measurement study
  • Redesigned thermalizing box with Paul
  • Created and posted BorAl shielded He3 attenuation measurements plot
  • Made and posted BorAl/Poly/Bpoly attenuation measurements plot


  • Fourth of July (observed)


  • Met with CoSE IT about DAQ issues
    • Need to upgrade DAQ to at least CENT OS 7, 8 may be better. Currently running version 6 which is too old for security updates
  • Worked with Paul on experimental setup design and marked some BorAl for cutting back at INL
  • Modeled experimental setup in fusion 360
    • Need more data from mcnp to finalize shielding designs
  • Drew cuts on BorAl for cutting


  • Washed regular poly for use in thermalizing cavity in experimental setup
  • Received new ADC and Gadolinium sheet from Chuting
  • Worked on MCNP and thermalizing box design with Paul
  • Gave some BorAl back to chuting to be cut


  • Downloaded new MCNP libraries to run Paul's new script
  • Spoke to Joe (chemist) about separating boron from BorAl and BPoly
  • Started working through python lessons
  • Started photon attenuation measurements in lead tunnel


  • Updated 3D model of feasibility study with Gadolinium converter and smoothed out some edges that were improperly situated
    • Next step is to model a cradle for He3 detector to be used in tunnel, then redesign for HPGe detector
  • Built experimental setup with Paul and attempted to get an energy spectrum through ADC


  • Compiled resources related to DAQ systems, specifically readout controller
  • Need to find a more reliable scope, may be one in LDS
  • Ran NaI through ADC and got energy spectrum. No calibration yet. Will update wiki later in week with fine details
    • Ran source out, source in, source in & Cadmium sheet
      • Noticeable changes in each


  • Adjusted thermalizing box design for easier source access
    • Took measurements of all poly needed
    • Preparing to cut poly to rebuild thermalizing box as designed
  • Designed cradle for He3 detector use in experimental setup
  • Continued python lessons
  • Emailed Josh Pak at ISU regarding chemical separations
  • Emailed Jack Bradley to get centos 7 disk
  • Did trial cuts on poly with Dr. Forest and found a method that gives accurate cuts


  • Continued fine tuning 3D design
  • Worked on more python lessons


  • Cut pieces of poly for thermalizing box
  • Built experimental setup
  • Discussed a run plan with Paul for next week
  • Created run plan for next week


  • Rendered images of experimental setup and uploaded to wiki page
  • Uploaded pictures taken of experimental setup
  • Finalized run plan for NaI and He3
  • Executed NaI run plan for preliminary measurements
  • Spoke to Roy Dunker about using Cs137, Co60, & Eu152 in clean room
  • Paperwork being filed with TSO


  • Executed He3 run plan for preliminary measurements
  • Met with Jeff Burggraf regarding ADC calibration and fixing NaI bases
  • Also spoke about using ADC and possible improvements


  • Ran both He3 and NaI with/without converter through ADC to create plots
  • Began disassembly of NaI bases to find issue
  • Both multimeters used were questionable, need 9V battery and fuses
  • Created plots of NaI and He3 measurements of past two days
  • Managed to copy files off of daq2 and onto minerve, cannot currently copy from minerve to local machine, had to rely on thumbstick
  • Cannot access root files easily without manual transfer before leaving ISU - will speak to CoSE IT about SCP issue
  • Have access to the source certificate for Cf252 source through Jeff Burggraf


  • Analyzed NaI and He3 energy spectrum data from ADC with Gd converter in/out
  • Presented data taken in experimental setup to INL collaborators
  • Discovered previously unknown artifact affecting data analysis on DAQ computer using various techniques including moving files to another machine for analysis and adjusting run parameters to give a known output. None were repeatable.


  • Continued to debug the DAQ system, most likely culprit is a bad RAM stick


  • Spoke with Tony regarding upgrading to new DAQ
  • Spoke with CoSE IT regarding upgrading to new DAQ
  • Started process of upgrading to new DAQ
  • First task is to get centos 7 up and running on a separate drive if possible. Currently the system cannot see the second disk through bios or centos 7 installer


  • Continued updating new DAQ
  • Mounted 1TB disk as storage for centos 7 under directory Data


  • Continued updating new DAQ, still


  • Continued updating new DAQ, even more
  • Planned rest of summer work with Paul regarding feasibility study



  • New DAQ has hit a wall with library issues not allowing evio2nt to parse dat files and create root files
  • Attempting solve issue on local machine instead of ISU property
  • Copied daq2 to minerve then to local machine
  • Ran power to work bench
  • Brought 3D printer to lab
  • HVAC updated research team, no AC in foreseeable future for clean room, so we get to continue to sweat
  • John Longley has given interim permission for sources to be used to calibrate HPGe in clean room
  • Attempted to build ROOT onto a new DAQ1 machine in hopes to skirt around issues with DAQ2


  • ROOT built successfully on Daq1
  • DAQ1 & DAQ2 can now analyze data
  • DAQ2 can also collect data


  • After a hiatus due to IT issues and a presentation, the HPGe detector is in the experimental setup

Future tasks not on meeting task list

Go Back